Find the cost per pound of a coffee mixture made from 8 lb of coffee that costs $9.20 per pound and 12 lb of coffee that costs $5.50 per pound.
Answer by kthanhpn(2)

The total amount of the new coffee mixture would be:
8 + 12 = 20 (pounds)
The price of 20 pounds of coffee mixture would be:
8(9.2) + 12(5.5) = 139.6 (dollars).
Therefore, the average price would be:
139.6/20=6.98 (dollars).
Answer: $6.98 per pound.